Weston Products manufactures an industrial cleaning compound that goes through three processing departments—Grinding, Mixing, and Cooking. All raw materials are introduced at the start of work in the Grinding Department. The Work in Process T-account for the Grinding Department for May is given below:

Work in Process—Grinding Department
Inventory, May 1 491,780 Completed and transferred
to the Mixing Department
?
Materials 859,040
Conversion 535,175
Inventory, May 31 ?

The May 1 work in process inventory consisted of 134,000 pounds with $300,160 in materials cost and $191,620 in conversion cost. The May 1 work in process inventory was 100% complete with respect to materials and 30% complete with respect to conversion. During May, 349,000 pounds were started into production. The May 31 inventory consisted of 47,000 pounds that were 100% complete with respect to materials and 70% complete with respect to conversion. The company uses the weighted-average method in its process costing system.

Required:

1. Compute the Grinding Department's equivalent units of production for materials and conversion in May.

2. Compute the Grinding Department's costs per equivalent unit for materials and conversion for May.

3. Compute the Grinding Department's cost of ending work in process inventory for materials, conversion, and in total for May.

4. Compute the Grinding Department's cost of units transferred out to the Mixing Department for materials, conversion, and in total for May.
Answer #1

1) Equivalent unit

EUP-Material EUP-Conversion
Unit transferred out 436000 436000
Ending WIP 47000 47000*70% = 32900
Total 483000 468900

2) Cost per equivalent unit

Material Conversion
Beginning WIP 300160 191620
Cost added 859040 535175
Total cost 1159200 726795
Equivalent unit 483000 468900
Cost per EUP 2.40 1.55

3) Cost of ending WIP

Material Conversion Total
Cost of ending WIP 47000*2.40 = 112800 32900*1.55 = 50995 163795

4) Cost of unit transferred out

Material Conversion Total
Cost of unit transferred out 436000*2.40 = 1046400 675800 1722200